OpenStudy (anonymous):

How do I complete conversions to the nearest hundredth? 10 cm = __ in.

OpenStudy (anonymous):

1 inch = 2.54 cm

OpenStudy (anonymous):

\[\frac{ 10 cm }{ ? } \times \frac{ 1 inch }{ 2.54 cm }\]

OpenStudy (anonymous):

Basically, \[\huge 10 \div 2.54=?\]

OpenStudy (anonymous):

3.93-->3.94

OpenStudy (anonymous):

Mathway_is the answer .254